You are here: Foswiki>Commissioning Web>Logs>31Mar2012_Log (02 Apr 2012, JuanGuerra)Edit Attach

AO Commissioning on 31 March 2012

Observer(s):C. Arcidiacono, L. Busoni, A.Riccardi,F. Quiros JC. Guerra (LBTO)

Telescope Operator: D. Huerta (LBTO)

SW Support: M delaPenia (Tucson)

Telescope Support: J. Urban (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

Close dome calibrations using the retroreflector.

The awake up of the left system. OK.

Power on the W#1. Ok. Checking the temperature trend of the BCU47.

Acquired the pupils and built a Reconstructor for BIN4. Check of the quality of the close loop at BIN3 and BIN4 using the laboratory table.

Installation of the neutral density filter.

Details

8:00 Power on the system, OK.

Board setup is saved to keep the actual W# cofiguration. Board setup saved: telescope-align6 Remove two actuators one from the current control and another one from the position. act #201 (current control) and #583 (position)

Hexapod position:

x(mm) y(mm) z(mm) Rx(arcsec) Ry(arcsec) Rz(arcsec)
4.6 1.5 4.58 317 612 0

During the Bin 3 test we can see the effect of an actuator on the flat saved. This actuator was #201 and it was removed from the current control.

Attuatore_2.png

BIN 3 calibration.

* Reconstructor with 10 modes and 66 modes. Test of the reconstructor at high flux with the disturbance was good. Check that the filter in the IRTC is J. * Test of the reconstructor with low flux (mag. 13) not yet, we need the density filter to get the proper flux on the ccd39.

20120330_212037 ??? 8.0 90 8 NaN -1.00 201828 3 66 800 0.8 0.6 0.0 6 4084 ON 14.8 J 90 30 ON NO - Warning: IRTC temperature > 260K
20120330_212116 ??? 8.0 90 9 NaN -1.00 201828 3 66 800 0.8 0.6 0.0 6 4083 ON 14.9 J 90 30 ON NO - Warning: IRTC temperature > 260K
20120330_212413 ??? 8.2 90 8 NaN -1.00 201828 3 66 800 0.8 0.6 0.0 6 3422 ON 15.1 J 90 30 ON NO - Warning: IRTC temperature > 260K

BIN 4 calibration.

* Reconstructor with 10 modes and 36 modes. Test of the reconstructor at high flux with the disturbance was good. Check that the filter in the IRTC is J. * Test of the reconstructor with low flux (mag. 1) not yet, we need the density filter to get the proper flux on the ccd39.

20120331_203107 ??? 8.4 90 4 NaN -1.00 200610 4 36 401 1.1 1.3 1.1 6 11952 ON 2.6 J 1000 10 ON NO  
20120331_203125 ??? 8.4 90 6 NaN -1.00 200610 4 36 401 1.1 1.3 1.1 6 11953 ON 2.5 J 1000 10 ON NO  
20120331_203604 ??? 8.3 90 4 NaN -1.00 200610 4 36 401 1.1 1.3 1.1 6 12665 ON 2.6 J 1000 10 ON NO  
20120331_203906 ??? 8.4 90 10 NaN -1.00 200610 4 36 401 1.7 1.2 0.0 6 12591 ON 2.0 J 1000 10 ON NO optimized gain manually
20120331_203936 ??? 8.3 90 10 NaN -1.00 200610 4 36 401 1.7 1.2 0.0 6 12597 ON 2.0 J 1000 10 ON NO optimized gain manually
20120331_204431 ??? 8.3 90 3 NaN -1.00 200610 4 36 401 1.1 1.1 1.3 6 12606 ON 2.5 J 1000 10 ON NO  
20120331_204504 ??? 8.3 90 9 NaN -1.00 200610 4 36 401 1.1 1.1 1.3 6 12603 ON 2.3 J 1000 10 ON NO  

Neutral density IN.

The installation and alignment of the neutral density filter at the W# unit took all the evening time. We couldn't perform any test after the installation.

-- JuanGuerra - 31 Mar 2012

  • act201:
    Attuatore_2.png
Topic revision: r4 - 02 Apr 2012, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback